40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 23 22:13:50 2010, Koji, Update, General, PZT HV turned on 
    Reply  Fri Dec 31 12:41:02 2010, rana, Update, General, IR flashes Untitled.png
Message ID: 4099     Entry time: Fri Dec 31 12:41:02 2010     In reply to: 4096
Author: rana 
Type: Update 
Category: General 
Subject: IR flashes 

I came in today to check up on the StripTool and burn the Ubuntu LTS (new CDS OS) DVD. I was pretty excited to see the PRM flashes on Mon1.

I waggled the PRM/BS alignments and got a good contrast MICH and then bright flashes in the PRM that totally overload Mon1's CCD.

Now I can see flashes of some IR junk in the X Arm; its way off on the left edge of the mirror, but there's a beam.

For the short term, we can hook up the IO PZTs to some old EPICS channels (like one of the AUX guys in the LSC area), but eventually it has to get hooked up to the new ASC or ASS. We have to bug Joe to see where this shows up in his master diagram.

**Note: if you get lost sometimes when doing the alignments, remember that you can use time_machine_conlog.

Ex:

rossa:general>./time_machine_conlog 2010/12/31,11:00:00 PDT C1:SUS-ITMX_PIT_COMM
Issuing the conlog command:
/cvs/cds/caltech/conlog/bin/conlog +epics -interp at 2010/12/31,11:00:00 PDT "C1:SUS-ITMX_PIT_COMM"


Conlog Replied:

LIGO controls: values at 2010 12/31 11:00:00 pst
 __Epics_Channel_Name______   __value_____
 C1:SUS-ITMX_PIT_COMM         0.406100



Restoring now...

Restoration Successful

Attached is the last 8 days of Vacuum Pressure trend which includes the pumpdown.

Attachment 1: Untitled.png  31 kB  Uploaded Fri Dec 31 13:51:49 2010  | Hide | Hide all
Untitled.png
ELOG V3.1.3-